OpenStudy (anonymous):

have a pitcher that holds 19 cups of water how many quarts is it

OpenStudy (anonymous):

1 quart = 4 cups. Use this to solve the problem.

OpenStudy (anonymous):

1/4=x/19 x=19*1/4 x=4.75 quarts Therefore the pitcher that holds 19 cups of water is 4.75 quarts.